suppose you are at a river resort and rent a motor boat for 5 hours starting at 7 am. You are told that the boat will travel at 8 mph upstream and 12 mph downstream. you decide that you want to go as far up the river as you can and still be back by noon. At what time should you turn back, and how far from the resort will you be at the time?
Answer by edjones(8007)

Let time it takes to go up the river be u and time it takes to go down be w.
u+w=5
u=5-w
.
d=st d=distance, s=speed, t=time
st=st since the distances are equal.
8u=12w
8(5-w)=12w substitute 5-w for u.
40-8w=12w
20w=40
w=2 hrs
u=3 hrs
8*2=16 mi distance at which you should turn back.
7 AM + 3 hrs= 10 AM time you should turn back.
.
Ed